OpenStudy (anonymous):

Evaluate: 32 + (7 − 2) • 4 − 6/3

OpenStudy (angela210793):

32+5*4-2=32+18=50

OpenStudy (anonymous):

calculate the brackets first

OpenStudy (anonymous):

50